OpenStudy (anonymous):

write the value missing from the perfect square trinomial x^2+ __ x +64

OpenStudy (anonymous):

Missing value=16

OpenStudy (campbell_st):

an expanded perfect square is (x + a)^2 = x^2 + 2ax + a^2 take the square root of 64 then double it to find your answer

OpenStudy (anonymous):

(x+8)(x+8)=x^2+8x +8x+64 x^2+16x+64